# Prediction-Powered Adaptive Shrinkage Estimation

### v1

@@ Appendix B Proofs of Theoretical Results
$\frac{4}{m^{2}}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\bigg{(}\sup_{\omega\geq 0}\bigg{|}\sum_{j=1}^{m}(1-\omega_{j})\big{[}\tilde{\gamma}_{j}-(\tilde{Z}^{f}_{j}-\mu_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\big{]}\bigg{|}\bigg{)}^{2}\right]\leq\frac{16}{m}\sup_{j}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\mathrm{Var}_{\eta_{j}}\!\left[(\tilde{Z}^{f}_{j}-\mu_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\right]\right],$
and so,
$\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\frac{2}{m}\sup_{\omega\geq 0}\bigg{|}\sum_{j=1}^{m}(1-\omega_{j})\big{[}\tilde{\gamma}_{j}-(\tilde{Z}^{f}_{j}-\mu_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\big{]}\bigg{|}\right]\leq\left(\frac{16}{m}\sup_{j}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\mathrm{Var}_{\eta_{j}}\!\left[(\tilde{Z}^{f}_{j}-\mu_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\right]\right]\right)^{1/2}.$
Again, the (fourth-)moment conditions from Lemma B.1 suffice to ensure that
@@
$\sup_{j}\mathbb{E}_{\mathbb{P}_{\eta}}\!\left[\mathrm{Var}_{\eta_{j}}\!\left[(\tilde{Z}^{f}_{j}-\mu_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\theta_{j})\right]\right]<\infty,$
and to establish control of
Step 4: Concluding the argument.
Finally, based on Steps 1–3, we have that
B.3 Proof of Theorem 5.2
Then:
Taking expectations,
B.4 Proof of Proposition 5.3
We start with the result in Theorem 5.2
Noting that $\tilde{\gamma}_{j}=0$ since $N_{j}=\infty$, we have
We finally have
Appendix C Other Baseline Shrinkage Estimators
C.1 “Shrink-classical” (Shrinkage) Baseline
@@ Appendix C Other Baseline Shrinkage Estimators
5:  $\hat{\omega}=\texttt{{get\_shrink\_param}}(\{\bar{Y}_{j}\}_{j=1}^{m},\{\tilde{Z}^{f}_{j}\}_{j=1}^{m},\{\tilde{\sigma}_{j}^{2}\}_{j=1}^{m})$triangleright use Eq. (C.1)
@@
7:    $\hat{\omega}_{j}\;=\;\hat{\omega}\Big{/}\bigl{(}\hat{\omega}+\tilde{\sigma}_{j}^{2}\bigr{)}$
@@
C.2 “Shrink-average” Baseline
$\bar{\theta}^{\mathrm{PT}}:=\frac{1}{m}\sum_{j=1}^{m}\hat{\theta}^{\mathrm{PT}}_{j}.$
Formulation.
The “shrink-average” estimator for problem $j$ takes the form:
$\hat{\theta}_{j,\omega}^{\text{Avg}}$ $:=\omega_{j}\hat{\theta}^{\mathrm{PT}}_{j}+(1-\omega_{j})\bar{\theta}^{\mathrm{PT}},$
where $\text{where }\quad\omega_{j}$ $:=\omega/(\omega+\tilde{\sigma}^{2}_{j}),\quad\tilde{\sigma}^{2}_{j}:=\mathrm{Var}_{\eta_{j}}[\hat{\theta}^{\mathrm{PT}}_{j}].$
Optimizing $\omega$ via SURE.
Xie et al. [2012] proposed the following unbiased risk estimate to optimize $\omega$ for this estimator. Note that even though the group mean is also correlated with each PT estimator, we still denote the following SURE instead of CURE following the nomenclature in Xie et al. [2012].
$\mathrm{SURE}(\hat{\theta}_{j,\omega}^{\text{Shrink}})$ $:=\left[(1-\omega_{j})(\hat{\theta}^{\mathrm{PT}}_{j}-\bar{\theta}^{\mathrm{PT}})\right]^{2}+(1-\omega_{j})(\omega+(2/m-1)\tilde{\sigma}_{j}^{2}).$
@@
0:  $(X_{ij},Y_{ij})_{i=1}^{n_{j}}$, $(\tilde{X}_{ij})_{i=1}^{N_{j}}$, $\rho_{j},\tau_{j},\sigma_{j}$ for $j\in[m]$, predictive model $f$
@@
5:    $\lambda_{j}^{*}=\texttt{{get\_pt\_param}}(\rho_{j},\tau_{j},n_{j},N_{j})$
@@
9:  $\bar{\theta}^{\mathrm{PT}}=m^{-1}\sum_{j=1}^{m}\hat{\theta}^{\mathrm{PT}}_{j}$
10:  $\triangleright$ Step 3: Adaptive shrinkage towards group mean (Eq. (30))
11:  $\hat{\omega}=\texttt{{get\_shrink\_param}}(\{\hat{\theta}_{j}^{\mathrm{PT}}\}_{j=1}^{m},\bar{\theta}^{\mathrm{PT}},\{\tilde{\sigma}_{j}^{2}\}_{j=1}^{m}$ )
@@
Appendix D Experiment Details
D.1 Sample-based Estimators for $\sigma_{j}^{2},\tau_{j}^{2},\gamma_{j}$
In 2.1, we assumed that the second-moment parameters $\sigma_{j}^{2},\tau_{j}^{2},\gamma_{j}$ at the sampling level are known in order to motivate our method and streamline theoretical results. In practice, when the true data generating process is unknown or when we only have access to a black-box predictor, we plug in the following sample-based (unbiased) estimators for the parameters:
$\hat{\sigma}_{j}^{2}$ $:=\frac{1}{n_{j}-1}\sum_{i=1}^{n_{j}}(Y_{ij}-\bar{Y}_{j})^{2},\quad\hat{\tau}_{j}^{2}:=\frac{1}{n_{j}+N_{j}-1}\bigg{(}\sum_{i=1}^{n_{j}}(f(X_{ij})-\hat{Z}_{j})^{2}+\sum_{i=1}^{N_{j}}(f(\tilde{X}_{ij})-\hat{Z}_{j})^{2}\bigg{)},$
These sample-based estimators are used in our numerical experiments involving real-world datasets.
D.2 Synthetic Model
@@
D.3 Amazon Review Ratings Dataset
@@
The Amazon Fine Food Reviews dataset, provided by the Stanford Network Analysis Project (SNAP; SNAP [2014]) on Kaggle, [^11] https://www.kaggle.com/datasets/snap/amazon-fine-food-reviews comes in a clean format. We group reviews by their ProductID. For each review, we concatenate the title and body text to form the covariate, while the response is the reviewer’s score/rating (1 to 5 stars). Here’s a sample review:
@@
The Bidirectional Encoder Representations from Transformers (BERT) model is a widely adopted language model for many NLP tasks including text classification [Devlin, 2018]. However, pretraining BERT from scratch is time-consuming and requires large amounts of data. We therefore use the bert-base-multilingual-uncased-sentiment model [^12] https://huggingface.co/nlptown/bert-base-multilingual-uncased-sentiment from NLP Town [2023] as the base model, denoted as BERT-base. BERT-base is pre-trained on general product reviews (not exclusive to Amazon) in six languages. It achieves 67.5% prediction accuracy on a validation set of 100 products ($\sim$46k reviews).
Then, we further fine-tune it on the held-out review data, that is, reviews outside the top 200 products, for 2 full epochs. The fine-tuning is done using Hugging Face’s transformers library [Wolf, 2019]. After fine-tuning, the BERT-tuned model achieves 78.8% accuracy on the same validation set.
D.4 Spiral Galaxy Fractions (Galaxy Zoo 2)
@@
The Galaxy Zoo 2 (GZ2) project [^13] https://data.galaxyzoo.org/ contains a large collection of human-annotated classification results for galaxy images from SDSS. However, instead of having a single dataframe, GZ2 has many different tables—each for subsets of the SDSS raw data. We begin with a particular subset of 239,696 images with metadata drawn from Hart et al. [2016]. Our data cleaning pipeline is inspired by Lin et al. [2021], which removes missing data and relabels the class name of each galaxy image to a more readable format:
@@
The additional SDSS metadata for GZ2141414The column names and their meanings are available at https://data.galaxyzoo.org/data/gz2/gz2sample.txt. contains valuable information that directly partitions the galaxies based on certain attributes, e.g., REDSHIFT_SIMPLE_BIN based on galaxy redshift measurements, and WVT_BIN calculated by weighted Voronoi tessellation. These partitions naturally motivate fine-grained compound mean estimation on this dataset.
After partitioning the images based on REDSHIFT_SIMPLE_BIN, [^15] In astronomy, REDSHIFT quantifies how much the wavelength of light from a galaxy is stretched due to the expansion of the universe. It serves as a good proxy for grouping galaxies of similar distance and time. we consider only the top $m=122$ partitions based on the cutoff that each problem should have $\geq 150$ images (many partitions have very few galaxy images in them), for the same reason as in the Amazon Review dataset. Finally, we have a total of $\sim$100k images as covariates (either $X_{ij}$ or $\tilde{X}_{ij}$) for our problem.
@@
We employ the ResNet50 architecture [He et al., 2016], utilizing the pre-trained model from torchvision initially trained on ImageNet [Deng et al., 2009]. To tailor the model to our task, we fine-tune it on $\sim$50k images excluded from the top $m$ problems. The model is trained to classify galaxies into eight categories, later condensed into a binary spiral/non-spiral classification for prediction. We use a batch size of 256 and Adam optimizer [Kingma, 2014] with a learning rate of 1e-3. After 20 epochs, the model achieves 87% training accuracy and 83% test accuracy. Despite these promising results, Table 3 indicates that the predictions still require de-biasing for accurate estimation.
D.5 Benchmarking in real-world datasets

### v2

@@ C.3 UniPAS: CURE and Adaptive Shrinkage with Estimated Second Moments
$\widehat{\mathrm{CURE}}\left(\boldsymbol{\hat{\theta}}^{\mathrm{UPAS}}_{\omega}\right)$ $:=\frac{1}{m}\sum_{j=1}^{m}\Big{[}(2\hat{\omega}_{j}-1)\dot{\sigma}^{2}_{j}+2(1-\hat{\omega}_{j})\dot{\gamma}_{j}+(1-\hat{\omega}_{j})^{2}\big{(}\hat{\theta}_{j}^{\mathrm{UPT}}-\tilde{Z}_{j}^{f}\big{)}^{2}\Big{]}.$ (26)
Finally, the UniPAS estimator is obtained by selecting the $\omega$ that minimizes this $\widehat{\mathrm{CURE}}$:
Definition C.4 (Univariate Prediction-Powered Adaptive Shrinkage (UniPAS)).
The UniPAS estimator for problem $j$ is $\hat{\theta}_{j}^{\mathrm{UPAS}}:=\hat{\theta}_{j,\hat{\omega}}^{\mathrm{UPAS}}$, where
$\hat{\omega}:=\arg\min_{\omega\geq 0}\widehat{\mathrm{CURE}}\left(\boldsymbol{\hat{\theta}}^{\mathrm{UPAS}}_{\omega}\right).$
This UniPAS estimator is fully data-driven and does not rely on knowledge of the true second-moment parameters. The pseudo-code for the full UniPAS algorithm is given below.
Algorithm 2 UniPAS
0:  $(X_{ij},Y_{ij})_{i=1}^{n_{j}}$, $(\tilde{X}_{ij})_{i=1}^{N_{j}}$ for $j\in[m]$, predictive model $f$
1:  for $j=1$ to $m$ do
2:     $\triangleright$ Step 1: Apply predictor (Eq. (1)) to get aggregated statistics and sample-based estimators for second moments
4:     $\hat{\sigma}_{j}^{2},\hat{\gamma}_{j},\hat{\tau}_{j}^{2}=\texttt{{get\_sample\_variances}}((X_{ij},Y_{ij})_{i=1}^{n_{j}},(\tilde{X}_{ij})_{i=1}^{N_{j}},f)$
5:     $\triangleright$ Step 2: Univariate power tuning (Section C.2)
6:     $\hat{\lambda}_{\mathrm{clip}}=\mathrm{clip}\bigg{(}\frac{\sum_{j=1}^{m}n_{j}^{-1}\hat{\gamma}_{j}}{\sum_{j=1}^{m}\frac{N_{j}+n_{j}}{N_{j}n_{j}}\hat{\tau}^{2}_{j}},[0,1]\bigg{)}$
7:     $\hat{\theta}_{j}^{\mathrm{UPT}}=\bar{Y}_{j}+\hat{\lambda}_{\mathrm{clip}}(\tilde{Z}^{f}_{j}-\bar{Z}_{j}^{f})$
8:     $\triangleright$ Step 3: Construct sample-based estimators of $\mathrm{Var}[\hat{\theta}_{j}^{\mathrm{UPT}}]$ and $\mathrm{Cov}[\hat{\theta}_{j}^{\mathrm{UPT}},\tilde{Z}_{j}^{f}]$
9:     $\dot{\sigma}_{j}^{2}=\frac{\hat{\sigma}^{2}_{j}}{n_{j}}+\frac{N_{j}+n_{j}}{N_{j}n_{j}}{\hat{\lambda}}_{\mathrm{clip}}^{2}\hat{\tau}_{j}^{2}-\frac{2}{n_{j}}\hat{\lambda}_{\mathrm{clip}}\hat{\gamma}_{j}$
10:     $\dot{\gamma}_{j}=\hat{\lambda}\frac{\hat{\tau}_{j}^{2}}{N_{j}}$
11:  end for
12:  $\triangleright$ Step 4: Adaptive shrinkage
13:  $\bar{\sigma}^{2}=\frac{1}{m}\sum_{j=1}^{m}\hat{\sigma}_{j}^{2},\>\bar{\tau}^{2}=\frac{1}{m}\sum_{j=1}^{m}\hat{\tau}_{j}^{2},\>\bar{\gamma}=\frac{1}{m}\sum_{j=1}^{m}\hat{\gamma}_{j}$
14:  for $j=1$ to $m$ do
15:     $\triangleright$ Calculate the averaging variance estimator in Equation 25
16:     $\check{\sigma}^{2}_{j}:=\frac{\bar{\sigma}^{2}}{n_{j}}+\frac{N_{j}+n_{j}}{N_{j}n_{j}}{\hat{\lambda}}_{\mathrm{clip}}^{2}\bar{\tau}^{2}-\frac{2}{n_{j}}\hat{\lambda}_{\mathrm{clip}}\bar{\gamma}$
17:  end for
18:  $\triangleright$ Now for each choice of $\omega\geq 0$, the shrinkage coefficient is determined by $\frac{\omega}{\omega+\check{\sigma}^{2}}$
19:  $\triangleright$ The CURE calculation in Equation 26, which still depends on the sample-based estimators $\dot{\sigma}_{j}^{2},\dot{\gamma}_{j}$
20:  $\hat{\omega}=\texttt{{get\_shrink\_param}}((\hat{\theta}_{j}^{\mathrm{UPT}})_{j=1}^{m},(\tilde{Z}^{f}_{j})_{j=1}^{m},(\check{\sigma}_{j}^{2})_{j=1}^{m},(\dot{\sigma}_{j}^{2})_{j=1}^{m},(\dot{\gamma}_{j})_{j=1}^{m})$
21:  for $j=1$ to $m$ do
22:     $\hat{\theta}_{j}^{\mathrm{UPAS}}=\hat{\omega}\hat{\theta}^{\mathrm{UPT}}_{j}+(1-\hat{\omega})\tilde{Z}^{f}_{j}$
23:  end for
24:  return$\>\{\hat{\theta}_{j}^{\textnormal{UPAS}}\}_{j=1}^{m}$
The fully data-driven construction of UniPAS is supported by the following theoretical guarantee:
Proposition C.5 (Asymptotic Consistency of $\widehat{\mathrm{CURE}}$ for UniPAS).
On top of the assumptions in Proposition C.2 and 2.2, if we further require that $\inf_{j\in[m],m\in\mathbb{N}}\mathring{\sigma}^{2}_{j,m}\geq\delta>0$ for some fixed $\delta$, where
$\mathring{\sigma}^{2}_{j}\equiv\mathring{\sigma}^{2}_{j,m}:=\frac{\mu_{\sigma^{2}}}{n_{j}}+\frac{N_{j}+n_{j}}{N_{j}n_{j}}(\lambda^{*}_{\mathrm{clip},m})^{2}\mu_{\tau^{2}}-\frac{2}{n_{j}}\lambda^{*}_{\mathrm{clip},m}\mu_{\gamma}$ (27)
with $\mu_{\sigma^{2}}:=\mathbb{E}_{\mathbb{P}_{\eta}}[\sigma_{j}^{2}],\,\mu_{\tau^{2}}:=\mathbb{E}_{\mathbb{P}_{\eta}}[\tau_{j}^{2}],\,\mu_{\gamma}:=\mathbb{E}_{\mathbb{P}_{\eta}}[\gamma_{j}]$. Then, $\widehat{\mathrm{CURE}}\left(\boldsymbol{\hat{\theta}}^{\mathrm{UPAS}}_{\omega}\right)$ is an asymptotically consistent estimator for the true loss $\ell_{m}(\boldsymbol{\hat{\theta}}^{\mathrm{UPAS}}_{\omega},\boldsymbol{\theta})$ of the UniPAS estimator (which uses weights $\hat{\omega}_{j}=\omega/(\omega+\check{\sigma}^{2}_{j})$), in the sense that:
@@
$\mathbb{E}_{\mathbb{P}_{\eta}}\left[\sup_{\omega\geq 0}|\widehat{\mathrm{CURE}}\left(\boldsymbol{\hat{\theta}}^{\mathrm{UPAS}}_{\omega}\right)-\ell_{m}(\boldsymbol{\hat{\theta}}^{\mathrm{UPAS}}_{\omega},\boldsymbol{\theta})|\right]\xrightarrow{m\to\infty}0.$
We defer the proof to Section F.7.
This consistency result ensures that minimizing $\widehat{\mathrm{CURE}}$ is asymptotically equivalent to minimizing the true MSE of the UniPAS estimator. This leads to the following optimality guarantee:
Theorem C.6 (Asymptotic Bayes Risk Optimality of UniPAS).
Under the assumptions of Theorem C.5, the UniPAS estimator $\hat{\boldsymbol{\theta}}^{\mathrm{UPAS}}=\hat{\boldsymbol{\theta}}^{\mathrm{UPAS}}_{\hat{\omega}}$ satisfies:
$\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\hat{\boldsymbol{\theta}}^{\mathrm{UPAS}}\Big{)}\leq\inf_{\omega\geq 0}\left\{\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\boldsymbol{\hat{\theta}}^{\mathrm{UPAS}}_{\omega}\Big{)}\right\}+o(1)\quad\text{ as }\;m\to\infty,$
and so
$\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\hat{\boldsymbol{\theta}}^{\mathrm{UPAS}}\Big{)}\leq\min\left\{\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\boldsymbol{\tilde{Z}^{f}}\Big{)},\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\hat{\boldsymbol{\theta}}^{\mathrm{UPT}}\Big{)}\right\}+o(1)\quad\text{ as }\;m\to\infty.$
The proof of Theorem C.6, which follows directly from Proposition C.5, mirrors the argument in in Section F.3 used to derive Theorem 5.2 from Proposition 5.1.
We note that the conclusion here is slightly weaker than that of Theorem 5.2 for PAS. Theorem 5.2 guarantees that PAS asymptotically has risk less or equal to that of PT with optimal per-problem choice of the power tuning parameter. By contrast, Theorem C.6 along with Theorem C.3 shows that UniPAS always has risk less or equal to that of power tuning that uses the same power tuning parameter for all problems. The main upshot of Theorem C.6 is that UniPAS does not require knowledge of second moments.
Appendix D Other Baseline Shrinkage Estimators
D.1 “Shrink-classical” (Shrinkage) Baseline
@@ Appendix D Other Baseline Shrinkage Estimators
5:  $\hat{\omega}=\texttt{{get\_shrink\_param}}((\bar{Y}_{j})_{j=1}^{m},(\tilde{Z}^{f}_{j})_{j=1}^{m},(\tilde{\sigma}_{j}^{2})_{j=1}^{m})$triangleright use Eq. (D.1)
@@
7:    $\hat{\omega}_{j}\;=\;\hat{\omega}/\bigl{(}\hat{\omega}+\tilde{\sigma}_{j}^{2}\bigr{)}$
@@
D.2 “Shrink-average” Baseline
@@
0:  $(X_{ij},Y_{ij})_{i=1}^{n_{j}}$, $(\tilde{X}_{ij})_{i=1}^{N_{j}}$, $\gamma_{j},\tau_{j},\sigma_{j}$ for $j\in[m]$, predictive model $f$
@@
5:    $\lambda_{j}^{*}=\texttt{{get\_pt\_param}}(\gamma_{j},\tau_{j},n_{j},N_{j})$
@@
9:  $\bar{\theta}^{\mathrm{PT}}=m^{-1}\sum_{j=1}^{m}\hat{\theta}^{\text{PT}}_{j}$
10:  $\triangleright$ Step 3: Adaptive shrinkage toward group mean (Eq. (29))
11:  $\hat{\omega}=\texttt{{get\_shrink\_param}}((\hat{\theta}_{j}^{\mathrm{PT}})_{j=1}^{m},\bar{\theta}^{\mathrm{PT}},(\tilde{\sigma}_{j}^{2})_{j=1}^{m})$
@@
$\bar{\theta}^{\mathrm{PT}}:=\frac{1}{m}\sum_{j=1}^{m}\hat{\theta}^{\text{PT}}_{j}.$
Formulation.
The “shrink-average” estimator for problem $j$ takes the form:
$\hat{\theta}_{j,\omega}^{\text{Avg}}$ $:=\omega_{j}\hat{\theta}^{\text{PT}}_{j}+(1-\omega_{j})\bar{\theta}^{\mathrm{PT}},$
where $\text{where }\quad\omega_{j}$ $:=\omega/(\omega+\tilde{\sigma}^{2}_{j}),\quad\tilde{\sigma}^{2}_{j}:=\mathrm{Var}_{\eta_{j}}[\hat{\theta}^{\text{PT}}_{j}].$
Optimizing $\omega$ via SURE.
Xie et al. (2012) proposed the following unbiased risk estimate to optimize $\omega$ for this estimator. Note that even though the group mean is also correlated with each PT estimator, we still denote the following SURE instead of CURE following the nomenclature in Xie et al. (2012).
$\mathrm{SURE}(\hat{\theta}_{j,\omega}^{\text{Shrink}})$ $:=\left[(1-\omega_{j})(\hat{\theta}^{\text{PT}}_{j}-\bar{\theta}^{\mathrm{PT}})\right]^{2}+(1-\omega_{j})(\omega+(2/m-1)\tilde{\sigma}_{j}^{2}).$
We refer to Algorithm 4 for a full pseudo-code implementation of “shrink-average” estimator.
Appendix E Experiment Details
E.1 Synthetic Model
@@
E.2 Amazon Review Ratings Dataset
@@
The Amazon Fine Food Reviews dataset, provided by the Stanford Network Analysis Project (SNAP; SNAP (2014)) on Kaggle, [^21] https://www.kaggle.com/datasets/snap/amazon-fine-food-reviews comes in a clean format. We group reviews by their ProductID. For each review, we concatenate the title and body text to form the covariate, while the response is the reviewer’s score/rating (1 to 5 stars). Here’s a sample review:
@@
The Bidirectional Encoder Representations from Transformers (BERT) model is a widely adopted language model for many NLP tasks including text classification (Devlin et al., 2019). However, pre-training BERT from scratch is time-consuming and requires large amounts of data. We therefore use the bert-base-multilingual-uncased-sentiment model [^22] https://huggingface.co/nlptown/bert-base-multilingual-uncased-sentiment from Town (2023) as the base model, denoted as BERT-base. BERT-base is pre-trained on general product reviews (not exclusive to Amazon) in six languages. It achieves 67.5% prediction accuracy on a validation set of 100 products ($\sim$46k reviews).
Then, we further fine-tune it on the held-out review data, that is, reviews outside the top 200 products, for 2 full epochs. The fine-tuning is done using Hugging Face’s transformers library (Wolf, 2019). After fine-tuning, the BERT-tuned model achieves 78.8% accuracy on the same validation set.
E.3 Spiral Galaxy Fractions (Galaxy Zoo 2)
@@
The Galaxy Zoo 2 (GZ2) project [^23] https://data.galaxyzoo.org/ contains a large collection of human-annotated classification results for galaxy images from SDSS. However, instead of having a single dataframe, GZ2 has many different tables—each for subsets of the SDSS raw data. We begin with a particular subset of 239,696 images with metadata drawn from Hart et al. (2016). Our data cleaning pipeline is inspired by Lin et al. (2021), which removes missing data and relabels the class name of each galaxy image to a more readable format:
@@
The additional SDSS metadata for GZ2242424The column names and their meanings are available at https://data.galaxyzoo.org/data/gz2/gz2sample.txt. contains valuable information that directly partitions the galaxies based on certain attributes, e.g., REDSHIFT_SIMPLE_BIN based on galaxy redshift measurements, and WVT_BIN calculated by weighted Voronoi tessellation. These partitions naturally motivate fine-grained compound mean estimation on this dataset.
After partitioning the images based on WVT_BIN, [^25] In the Galaxy Zoo 2 dataset, WVT_BIN denotes Voronoi bins constructed based on each galaxy’s intrinsic size and absolute magnitude. The motivation and implementation of this binning strategy are detailed in Hart et al. (2016), who justify such partitioning—aligned with our compound mean estimation setup—by noting that spiral arm morphology exhibits systematic dependencies on stellar mass and related intrinsic properties. we consider only the top $m=100$ partitions based on the cutoff that each problem should have $\geq 150$ images (many partitions have very few galaxy images in them), for the same reason as in the Amazon Review dataset. Finally, we have a total of $\sim$100k images as covariates (either $X_{ij}$ or $\tilde{X}_{ij}$) for our problem.
@@
We employ the ResNet50 architecture (He et al., 2016), utilizing the pre-trained model from torchvision initially trained on ImageNet (Deng et al., 2009). To tailor the model to our task, we fine-tune it on $\sim$50k images excluded from the top $m$ problems. The model is trained to classify galaxies into eight categories, later condensed into a binary spiral/non-spiral classification for prediction. We use a batch size of 256 and Adam optimizer (Kingma & Ba, 2015) with a learning rate of 1e-3. After 20 epochs, the model achieves 87% training accuracy and 83% test accuracy. Despite these promising results, Table 3 indicates that the predictions still require debiasing for accurate estimation.
E.4 Benchmarking in real-world datasets